Topics in Reasoning

April 10, 1996

Sunil Logics for monotonic Reasoning

– Propositional logic - syntax, semantics, validity, models, inference rules and complexity of propositional inference

– 1st-Order logic

∗ Syntax,Semantics,Extensions, Notations, Axioms, Definitions, Theorems

∗ Representing Change (e.g. Situation Calc.)

∗ Inference (Generalized Modus Ponens, Forward and Backward

Chaining, Completeness, Resolution)

– High order logics

• Nonmonotonic Reasoning Urszula Default Logic Urszula Autoepistemic Logic Urszula KLM properties Urszula κ ranking Urszula ϵ semantics Urszula Modal Logics - Belief and Knowledge

Eyal Circumscription

Eyal Abduction

Eyal A Robot learning using Circumscription (Shanahan AAAI 96’)

Eyal Inheritance

Eyal Modal Logics beyond Belief and Knowledge

Eyal The Nonmonotonic Inference Relation |∼ beyond KLM. ? Temporal Reasoning

Pederito Logical Reasoning Systems

– Theorem Provers (SAM, AURA, OTTER, PVS) and Logic Programming (Prolog, MRS, LIFE).

– Production Systems(OPS5, CLIPS, SOAR).

– Frame Systems (OWL, FRAIL, KODIAK) and Semantic Networks (SNEPS, NETL, Conceptual Graphs).

– Description Logic systems (KL1, CLASSICS, LOOM). Urszula TMSs, CMS.

Avi? Probabilistic Reasoning – axioms,rules and the probabilities origins.

– Belief Networks

∗ Buz-words: Canonical Distribution, Deterministic Nodes, Noisy-OR, Leak Node, D-separation, Blocked.

∗ Inference in Belief Networks

∗ Buz-words: Diagnostic Inference, Causal Inference, Intercausal Inference, Mixed Inference.

∗ Clustering Methods for BNs

∗ Knowledge Engineering

– Decision Networks and Utility

– Sequential Decision Problems

– Rule-based methods, Certainty Factors and MYCIN

– The Dempster-Shafer theory.

– Qualitative Probabilistic Networks

– Examples: HUGIN, CONVINCE, MUNIN, PATHFINDER

• Fuzzy sets and Fuzzy logic

• Deductive and Inductive Reasoning.

Topic to deal with: Planning vs Reasoning

This is /u/eyal/ongo/qual/reasoning topics.tex.